I need to come up with an computation where i split $100,000. I need for each boy (b) to make 75% more than each girl (g). If i have 65 boys and 34 girls what would the formula look like so that i could put the numbers in so that this works out?
Answer by ankor@dixie-net.com(22740)

I need to come up with an computation where i split $100,000.
I need for each boy (b) to make 75% more than each girl (g).
If i have 65 boys and 34 girls what would the formula look like so that i could put the numbers in so that this works out?
:
That means
b = 1.75g
:
65b + 34g = 100000
replace b with 1.75g
65(1.75g) + 34g = 100000
113.5g + 34g = 100000
147.75g = 100000
g = 10000/147.75
g = $676.82 for each girl
and
1.75(676.82) = $1184.43 for each boy